# A Catalog of $\sim 146,000$ Radio Galaxies up to $z\sim 5$

### v1

@@ 2.1 UNIONS
Figure 2: The angular separation-$\mathtt{NN\_dist}$ plane for UNIONS-VLASS sources cross-matched with a search radius of $6^{\prime\prime}$. The vertical lines indicate the $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ and the horizontal line represents search radius to estimate completeness and contamination.
2.2 VLASS
The VLASS (Very Large Array Sky Survey; Lacy et al. 2020) observes the entire sky at a declination $>-40$ degrees, covering a total area of 33,885 $\mathrm{deg^{2}}$. The observation frequency ranges $2-4$ GHz, with a central (reference) frequency of 3 GHz. In baseline configurations in B and BnA to compensate for the projection effect, an angular resolution of $2.5\arcsec$ has been reached. The total survey is divided into three epochs, which have been completed in October 2024, plus a further half-epoch observation. [^1] https://science.nrao.edu/vlass/library/vlass-epoch-4-science-case Plus the single-epoch rms sensitivity (root mean square error $1\sigma$) that reaches 120 $\mu$Jy, these features make VLASS the widest and most sensitive radio sky survey with an excellent resolution at 3 GHz (de Gasperin et al., 2021). Therefore, VLASS is ideal for serving as our reference radio source catalog.
VLASS data release delivers quick look (QL) and single epoch (SE) images. QL images are products of the rapid-CLEAN process with a cell size of $1\arcsec$. SE images undergo deeper cleaning and self-calibration and reach a better cell size of $0.6\arcsec$. However, only an area of 1,000 $\mathrm{deg^{2}}$ is available in SE, we therefore only use QL products for the cross-match. We use the epoch 2 catalog (VLASS2 hereafter) since astrometric errors and underestimates of flux densities were previously present in epoch 1, and the final release of epoch 3 is unsettled.
There are $\sim 3$ M radio sources detected in VLASS2. Prior to cross-matching, we adopted the following criteria to filter the VLASS2 catalog following the VLASS catalog user guide (see also Gordon et al. 2021)
@@ 2.2 VLASS
  • $\mathtt{S\_code!=E}$: the source can be fitted with a single-Gaussian or multi-Gaussian components;
plus $\mathtt{NN\_dist}$, which the angular separation between the target source and its nearest neighbor in VLASS2, is larger than $30^{\prime\prime}$ (see § 3.1 for the discussion). These criteria reduce the number of sources in “clean” VLASS2 to 1,380,707. The VLASS2 catalog also includes the column $\mathtt{P\_sidelobe}$, which means the probability of a source being a sidelobe but misclassified as a real source. As recommended in the VLASS Catalog User Guide, $\mathtt{Quality\_flag==0}$ that indicates $\mathtt{P\_sidelobe<0.1}$ is preferred, and the estimated contamination rate is 0.39%. In the cleanest criterion, $\mathtt{P\_sidelobe<0.05}$ is applied, and the estimated contamination rate is 0.12%. For the final “clean” sample of 146,212 UNIONS-VLASS sources cross-matched (see § 3.3 for details), adopting these two criteria results in a difference of 1,766 ($\sim 1.2\%$) sources, which is statistically insignificant. In the following of this work, without the loss of totality of our catalog, we present results adopting $\mathtt{P\_sidelobe<0.1}$.
@@ 2.3 FIRST
FIRST (Becker et al., 1995) employs VLA to observe 8,444 $\mathrm{deg^{2}}$ square degrees of the NGC at 1.4 GHz and the survey area is limited to a declination of $<65$ degrees. In the SGC, an area of $644\mathrm{\ deg^{2}}$ has been observed, whereas the survey area does not overlap with the current internal data release of UNIONS5000 (see Figure 5). In this work, we use the FIRST final data release catalog (Helfand et al., 2015). In B-configuration, the survey reaches a spatial resolution represented by a circular Gaussian with full width at half maximum (FWHM) $5.4\arcsec$ with a pixel size of $1.8\arcsec$ over the northern sky. For bright, pointlike sources with measured extents of $\leq 2.0\arcsec$, the statistical astrometry is $0.22\arcsec$. In a typical situation, the flux density detection limit is 1 mJy over the full footprint.
@@ 2.4 LoTSS
To understand the evolution of radio galaxies, the spectral shape plays a crucial role, which requires multifrequency observations to construct the synchrotron spectral energy distribution (SED). MHz-observations carried out by LOw-Frequency ARray are ideal supplements to our GHz-VLASS and FIRST data. The LoTSS (LOFAR Two-metre Sky Survey; Shimwell et al. 2017, 2022) employs the LOFAR to observe 27% of the northern sky at $120-168$ MHz, with a central (reference) frequency of 144 MHz. The sky coverage has two separate regions that span 4,178 and 1,457 $\mathrm{deg^{2}}$ respectively. The full bandwidth Stokes I continuum imaging of LoTSS yields images at a spatial resolution of $6.0\arcsec$, centered at a frequency of 144 MHz. The median detection sensitivity of the survey is 89 $\mathrm{\mu Jy\ beam^{-1}}$ with a typical positional uncertainty smaller than $0.2\arcsec$. In this work, we use the latest and widest coverage catalog, LoTSS DR2 (Shimwell et al., 2022), which contains 4,396,228 radio sources derived from integrated intensity maps in Stokes I polarization.
@@ 3.1 $\mathtt{NN\_dist}$ Cut
Radio galaxies with low radio luminosities often have their jets confined by the ISM and rapidly disrupted (Fanaroff & Riley, 1974; Turner & Shabala, 2015). As a result, the peak surface brightness can be close to the host galaxy. The nearest neighbour matching is thus a straightforward and efficient approach for identifying the optical counterpart of a radio source (Williams et al., 2019). On the other hand, powerful radio jets can rapidly overcome the confinement of the ISM of their host galaxies without suffering significant disruption (Mukherjee et al., 2016). Thereby, they often propagate far from their host galaxies and end in the terminal shocks known as radio hotspots, where the observed radio surface brightness peaks (Turner & Shabala, 2023). Identifying the optical counterparts of such radio sources requires a larger search radius, which, in turn, inevitably increases the likelihood of contamination by unrelated sources (Banfield et al., 2015; Gordon et al., 2023).
@@
To statistically avoid erroneous associations attributed to large-scale, complex radio structures, we perform a cross-match between UNIONS and the VLASS2 catalog – adopting the criteria present in Gordon et al. (2021) – choosing a search radius of $6^{\prime\prime}$, which is $\sim 550,000$ sources. The angular separation between UNIONS-VLASS versus $\mathtt{NN\_dist}$ is shown in Figure 2. It is clear that the number of cross-matched sources shows a bimodal distribution along $\mathtt{NN\_dist}$. Sources cross-matched with $\mathtt{NN\_dist<30^{\prime\prime}}$ tend to have large angular separations between UNIONS and VLASS centroids. We randomly select 36 VLASS sources with $\mathtt{NN\_dist<10^{\prime\prime}}$ and show their QL images, overlaid with VLASS and UNIONS source centroids, in Figure 18. For these radio sources, they often show complex and extended radio morphology, including core-jet and double-jet structures. VLASS could capture the centroid of the lobe or radio hotspot, naturally begetting the large offsets between UNIONS and VLASS. We, therefore, choose a $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ to avoid false positive associations between UNIONS-VLASS arising from evolved radio structures that can extend well beyond host galaxy scale and overlap with other galaxies. This cut can remove a large fraction of bright, evolved radio galaxies at low-$z$. Consequently, we see that UNIONS-VLASS RGs are absent from the bright end (${F}_{\mathrm{3GHz}}>3000$ mJy; Figure 9) and could be biased toward compact, young radio AGN populations (Figures 1516; see § 5.3 for discussion).
@@ 3.2 Contamination and Completeness
Without excluding optical counterparts associated with a single radio source, we show the differential distribution of the separation between UNIONS and VLASS sources (rightmost panel of Figure 3). At ${r}_{\mathrm{s}}>1.5\arcsec$, the increase in the search radius is accompanied by an increase in the number of optical counterpart counts in the annulus, suggesting these optical counterparts are likely to be coincident associations rather than real matches. The threshold of ${r}_{\mathrm{s}}\sim 1.5\arcsec$ is consistent with the completeness-cleanness trade-off of $1.6\arcsec$ adopting a median galaxy number density of $36.1$ per $\mathrm{arcmin^{2}}$. This value is also comparable to Gordon et al. (2023), who found that the probability that an AllWISE soruce is the true positive counterpart of an unresolved VLASS radio source reaches 80% at a positional uncertainty of $1.8\arcsec$.
@@ 3.3 Clean UNIONS-VLASS Radio Galaxies
If the radio source has more than one optical counterpart, we always choose the nearest one as the real match. This yields 10,541 radio sources that have multiple optical sources within a search radius of $1.5\arcsec$, and they are flagged with $\mathtt{1}$ in the column $\mathtt{flag\_multi\_opt}$. Because of the general trend of being compact for our UNIONS-VLASS RGs (see § 4.1 and § 5.1), this nearest match method could be valid in most cases, as jets and jet-induced shocks have not traveled far away from the host galaxy. However, it should be noted that this does not fully rule out optical counterparts by chance overlaps. Especially in cases where the RGs have developed jet-lobe structures that can result in large offsets between radio-emitting regions and host galaxies. Consequently, the centroid of the radio emission captured by VLASS may be the radio hotspot and appears to be closer to another galaxy (up to approximately $4\%$ based on $\sim 1,500$ Euclid-VLASS RGs; Zhong et al. in prep.). These multiple optical counterparts are also considered merging or interacting galaxy candidates, and they will be investigated in detail in future work.
The results of the cross-match are summarized in Table 1. Our angular separation-based cross-match generated 146,212 VLASS radio sources matched to UNIONS optical counterparts. Defining the fraction of matched VLASS sources in the area that overlaps with the UNIONS as the fraction of matches $({f}_{\mathrm{matches}})$, the corresponding value is 0.594, which is the maximum to be considered. By virtue of the higher sensitivity and wide sky coverage provided by the UNIONS, this rate is higher than those in the previous studies cross-matching FIRST with SDSS ($\sim 30\%$; Ivezić et al. 2002; Helfand et al. 2015). This fraction of matches also exceeds the cross-matching FIRST with Subaru HSC ($\sim 47\%$; Yamashita et al. 2018), even HSC optical number counts share a similar trend as the UNIONS (see 4.2 for further discussions). This could be explained by a combination of several factors, including a higher sensitivity of VLASS over FIRST and a larger search radius ($1.5\arcsec$) of UNIONS-VLASS than $1\arcsec$ of HSC-FIRST.
@@ 3.4 UNIONS-VLASS Astrometry
This vast radio galaxy catalog offers an excellent opportunity to investigate the systematic astrometric offsets between UNIONS and VLASS catalogs. The offsets along R.A. and DEC. are residuals of a VLASS position minus a UNIONS position: $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$ and $\Delta\delta=\mathrm{\delta_{VLASS}-\delta_{UNIONS}}$. However, our cross-match method selects associated pairs based on the angular separation ($\Delta\theta$) between radio and optical sources. Because the positional difference along the R.A. on the celestial sphere has a declination dependence, this calculation has to be corrected for the projection effect and is given by $\Delta\theta=\sqrt{({\Delta\alpha}_{\mathrm{corr}})^{2}+(\Delta\delta)^{2}}$, ${\Delta\alpha}_{\mathrm{corr}}=(\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}})\times\cos({\delta}_{\mathrm{UNIONS}})$. Therefore, for $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$, the value can exceed the search radius of $1.5\arcsec$, and such an excess becomes more severe for sources at high declinations, as shown in the left panel of Figure 19.
@@
The standard deviations of 140,065 UNIONS-VLASS sources whose $\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}\leq 1.5\arcsec$ are $0.44\arcsec$ in R.A. and $0.29\arcsec$ in DEC. On the other hand, for 146,212 “clean” UNIONS-VLASS sources, the offset along the R.A. retards to $0.61\arcsec$ while the change in the offset along DEC. is negligible. These offsets are much worse than the positional uncertainties (20 mas) of the CFHT $u-$ and $r-$band imaging surveys employed by UNIONS (Gwyn et al., 2025). For the VLASS catalog, by comparing the Epoch 1 catalog with Gaia DR2, the VLASS component positional accuracy is $\sim 0.5\arcsec$ at $\mathrm{DEC.>-20\ deg^{2}}$. Therefore, the current standard deviations of the positional errors for our UNIONS-VLASS sources are thus reasonable compared to the entire VLASS survey, suggesting that the astrometric offsets between optical and radio pairs are attributed to the intrinsic uncertainties of VLASS QL images.
In VLASS Epoch 1, the positional error peaks at $0.0\arcsec$ in R.A. and peaks at $-0.25\arcsec$ in DEC. Our UNIONS-VLASS sources based on Epoch 2 observations show a similar trend in R.A. with an improved astrometric accuracy in DEC. (${\Delta\delta}_{\mathrm{peak}}\sim 0.0\arcsec$) compared to VLASS Epoch 1. However, there is no significant improvement in the typical astrometry of $\sim 0.5\arcsec$, which is much worse than that of FIRST ($0.2\arcsec$). This is more of a problem of the rapid CLEAN procedure adopted by VLASS QL products. More accurate astrometry can be expected when Epoch 3 and epoch-concatenated data are delivered, as well as for the SE images.

### v2

@@ 2.2 VLASS
VLASS data releases deliver quick look (QL) and single epoch (SE) images. QL images are products of the rapid-CLEAN process with a cell size of $1\arcsec$. SE images undergo deeper cleaning and self-calibration and reach a better cell size of $0.6\arcsec$. However, since only an area of 1,000 $\mathrm{deg^{2}}$ is available in the SE images, we exclusively use QL products for the cross-match. We use the epoch 2 catalog (VLASS2 hereafter) since astrometric errors and underestimates of flux densities were previously present in epoch 1, and the final release of epoch 3 is unsettled.
There are $\sim 3$ M radio sources detected in VLASS2. Prior to the cross-match, we adopt the following criteria to filter the VLASS2 catalog following the VLASS catalog user guide (see also gordon_quick_2021)
@@
  • $\mathtt{S\_code!=E}$: the source can be fitted with a single-Gaussian or multi-Gaussian components.
We further apply a cut on $\mathtt{NN\_dist}$ – the angular separation between a VLASS source and its nearest neighbor in VLASS2 – requiring it to be greater than $30^{\prime\prime}$ (see § 3.1 for details). These criteria reduce the number of sources in “clean” VLASS2 to 1,380,707. The VLASS2 catalog also includes the column $\mathtt{P\_sidelobe}$, which means the probability of a source being a sidelobe but misclassified as a real source. As recommended in the VLASS Catalog User Guide, $\mathtt{Quality\_flag==0}$ that indicates $\mathtt{P\_sidelobe<0.1}$ is preferred, and the estimated contamination rate is 0.39%. In the cleanest criterion, $\mathtt{P\_sidelobe<0.05}$ is applied, and the estimated contamination rate is 0.12%. For the final “clean” sample of 146,212 UNIONS-VLASS sources cross-matched (see § 3.3 for details), adopting these two criteria results in a difference of 1,766 ($\sim 1.2\%$) sources, which is statistically insignificant. In the following of this work, without the loss of totality of our catalog, we present results adopting $\mathtt{P\_sidelobe<0.1}$.
@@ 2.3 FIRST
FIRST (becker_first_1995) employs VLA to observe 8,444 $\mathrm{deg^{2}}$ square degrees of the NGC at 1.4 GHz and the survey area is limited to a declination of <65<65 degrees. In the SGC, an area of $644\mathrm{\ deg^{2}}$ has been observed, whereas the survey area does not overlap with the current internal data release of UNIONS5000 (see Figure 5). In this work, we use the FIRST final data release catalog (helfand_last_2015). In B-configuration, the survey reaches a spatial resolution represented by a circular Gaussian with full width at half maximum (FWHM) $5.4\arcsec$ with a pixel size of $1.8\arcsec$ over the northern sky. For bright, pointlike sources with measured extents of $\leq 2.0\arcsec$, the statistical astrometry is $0.22\arcsec$. In a typical situation, the flux density detection limit is 1 mJy over the full footprint.
@@ 2.4 LoTSS
To understand the evolution of radio galaxies, the spectral shape plays a crucial role, which requires multifrequency observations to construct the synchrotron spectral energy distribution (SED). MHz-observations carried out by LOw-Frequency ARray are ideal supplements to our GHz-VLASS and FIRST data. The LoTSS (LOFAR Two-metre Sky Survey; shimwell_lofar_2017; shimwell_lofar_2022) employs the LOFAR to observe 27% of the northern sky at 120168120-168 MHz, with a central (reference) frequency of 144 MHz. The sky coverage has two separate regions that span 4,178 and 1,457, $\mathrm{deg^{2}}$ respectively. The full bandwidth Stokes I continuum imaging of LoTSS yields images at a spatial resolution of $6.0\arcsec$, centered at a frequency of 144 MHz. The median detection sensitivity of the survey is 89 $\mathrm{\mu Jy\ beam^{-1}}$ with a typical positional uncertainty smaller than $0.2\arcsec$. In this work, we use the latest and widest coverage catalog, LoTSS DR2 (shimwell_lofar_2022), which contains 4,396,228 radio sources derived from integrated intensity maps in Stokes I polarization.
@@ 3.1 $\mathtt{NN\_dist}$ Cut
Radio galaxies with low radio luminosities often have their jets confined by the ISM and rapidly disrupted (fanaroff_morphology_1974; turner_energetics_2015). Resultantly, the position of the peak surface brightness can be close to the host galaxy. The nearest neighbor cross-match is thus a straightforward and efficient approach for identifying the optical counterpart of a radio source (williams_lofar_2019). On the other hand, powerful radio jets can quickly overcome the confinement imposed by the ISM of their host galaxies without experiencing significant disruption (mukherjee_relativistic_2016). Consequently, they often propagate far from their host galaxies and terminate in strong shocks – referred to as radio hotspots – where the observed radio surface brightness typically peaks (turner_dynamics_2023). Identifying the optical counterparts of such radio sources necessitates a larger search radius, which, in turn, inevitably brings about more contaminants (banfield_radio_2015; gordon_quick_2023).
@@
To statistically minimize chance overlaps between optical sources and radio sources exhibiting large-scale, complex structures, we first perform a cross-match between UNIONS and VLASS2 catalogs adopting the criteria present in gordon_quick_2021 and choosing a search radius of $6^{\prime\prime}$. This cross-match yields $\sim 550,000$ sources. The angular separation between a pair of UNIONS-VLASS sources versus $\mathtt{NN\_dist}$ of the corresponding VLASS source is shown in Figure 2. It is clear that the number of sources that are cross-matched shows a bimodal distribution along $\mathtt{NN\_dist}$. Cross-matched UNIONS-VLASS pairs with $\mathtt{NN\_dist\leq 30^{\prime\prime}}$ tend to have large angular separations between their UNIONS and VLASS centroids. We randomly select 36 VLASS sources with $\mathtt{NN\_dist<10^{\prime\prime}}$ and show their QL images, overlaid with VLASS and UNIONS source centroids, in Figure 18. These radio sources often exhibit complex and extended morphology, such as core-jet and double-jet structures. VLASS could capture the centroid of the lobe or radio hotspot, naturally leading to the large offsets between UNIONS and VLASS. We, hence, choose a $\mathtt{NN\_dist}$ cut of $30^{\prime\prime}$ to avoid false positive associations between UNIONS-VLASS caused by evolved radio structures that can extend well beyond the host galaxy scale and overlap with unrelated galaxies.
It should be noted that this cut can remove a large fraction of bright, evolved radio galaxies at low-$z$. Quantitatively, for a given search radius, let M denote the number of cross-matched sources with $\mathtt{NN\_dist\leq 30^{\prime\prime}}$ and N the number with $\mathtt{NN\_dist>30^{\prime\prime}}$. The fractional loss of the completeness can then be expressed as ${f}_{\mathrm{loss}}=M/(N+M)$. As the search radius decreases from $6^{\prime\prime}$ to $1.5^{\prime\prime}$, ${f}_{\mathrm{loss}}$ decreases from $\sim 25\%$ – roughly the fraction of sources with $\mathtt{NN\_dist\leq 30^{\prime\prime}}$ in the VLASS2 catalog – to approximately $12\%$, while the number drops from 64,275 to 19,539. Therefore, in addition to $\mathtt{NN\_dist}$, the elimination of extended radio sources with complex structures is also sensitive to the search radius resulting from the design of the cross-match methodology described in the following sections. Consequently, we see that UNIONS-VLASS RGs are absent from the bright end (${F}_{\mathrm{3GHz}}>3000$ mJy; Figure 9) and could be biased toward compact, young radio AGN populations (Figures 1516; see § 5.3 for details).
@@ 3.2 Contamination and Completeness
Without excluding optical counterparts associated with a single radio source, we show the differential distribution of the separation between UNIONS and VLASS sources (rightmost panel of Figure 3). At ${r}_{\mathrm{s}}>1.5\arcsec$, the increase in the search radius is accompanied by an increase in the number of optical counterpart counts in the annulus, suggesting that these optical counterparts are likely to be coincident associations rather than real matches. The threshold of ${r}_{\mathrm{s}}\sim 1.5\arcsec$ is consistent with the completeness-cleanness trade-off of $1.6\arcsec$ adopting a median galaxy number density of $36.1$ per $\mathrm{arcmin^{2}}$. This value is comparable to the findings of gordon_quick_2023, who reported that the probability that an AllWISE source is the true postive counterpart to an unresolved VLASS radio source reaches 80% at a positional uncertainty of $1.8^{\prime\prime}$.
@@ 3.3 Clean UNIONS-VLASS Radio Galaxies
If a radio source has multiple optical counterpart candidates, we select the nearest one as the most likely true match. This yields 10,541 radio sources with multiple optical counterparts within a search radius of $1.5\arcsec$, which are flagged with a value of $\mathtt{1}$ in the column $\mathtt{flag\_multi\_opt}$. These multiple optical counterparts are also considered merging or interacting galaxy candidates, and they will be investigated in detail in future work. Given the general compactness of our UNIONS-VLASS RGs (see § 4.1 and § 5.1), this nearest match method is expected to be valid in most cases, as jets and jet-induced shocks have not propagated far from their host galaxies. However, this methodology does not fully rule out coincidental associations with unrelated optical counterparts, particularly in cases where RGs have developed jet-lobe structures that can lead to large offsets between radio-emitting regions and host galaxies. As a result, the centroid of the radio emission captured by VLASS may be the radio hotspot and appears to be closer to an unassociated galaxy (up to approximately $6\%$ based on $\sim 1,000$ Euclid-VLASS RGs; Zhong et al. in prep.).
The results of the cross-match are summarized in Table 1. Our angular separation-based cross-match generated 146,212 VLASS radio sources matched to UNIONS optical counterparts. Defining the fraction of matched VLASS sources in the area that overlaps with the UNIONS as the fraction of matches $({f}_{\mathrm{matches}})$, the corresponding value is 0.594, which is the maximum to be considered. By virtue of the higher sensitivity and wide sky coverage provided by the UNIONS, this rate is higher than those in the previous studies cross-matching FIRST with SDSS ($\sim 30\%$; ivezic_optical_2002; helfand_last_2015). This fraction of matches also exceeds the cross-matching FIRST with Subaru HSC ($\sim 47\%$; yamashita_wide_2018), even HSC optical number counts share a similar trend as the UNIONS (see 4.2 for further discussions). This could be explained by a combination of several factors, including a higher sensitivity of VLASS over FIRST and a larger search radius ($1.5\arcsec$) of UNIONS-VLASS than $1\arcsec$ of HSC-FIRST.
@@ 3.4 UNIONS-VLASS Astrometry
This vast radio galaxy catalog offers an excellent opportunity to investigate the systematic astrometric offsets between UNIONS and VLASS catalogs. The offsets along R.A. and DEC. are residuals of a VLASS position minus a UNIONS position: $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$ and $\Delta\delta=\mathrm{\delta_{VLASS}-\delta_{UNIONS}}$. However, our cross-match method selects associated pairs based on the angular separation ($\Delta\theta$) between radio and optical sources. Because the positional difference along the R.A. on the celestial sphere has a declination dependence, this calculation has to be corrected for the projection effect and is given by $\Delta\theta=\sqrt{({\Delta\alpha}_{\mathrm{corr}})^{2}+(\Delta\delta)^{2}}$, where ${\Delta\alpha}_{\mathrm{corr}}=(\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}})\times\cos({\delta}_{\mathrm{UNIONS}})$. Therefore, for $\Delta\alpha=\mathrm{\alpha_{VLASS}-\alpha_{UNIONS}}$, the value can exceed the search radius of $1.5\arcsec$, and such an excess becomes more severe for sources at higher declinations, as shown in the left panel of Figure 19.
@@
The standard deviations of 140,065 UNIONS-VLASS sources whose $\sqrt{\Delta\alpha^{2}+\Delta\delta^{2}}\leq 1.5\arcsec$ are $0.44\arcsec$ in R.A. and $0.29\arcsec$ in DEC. On the other hand, for 146,212 “clean” UNIONS-VLASS sources, the offset along the R.A. retards to $0.61\arcsec$ while the change in the offset along DEC. is negligible. These offsets are much worse than the positional uncertainties (20 mas) of the CFHT $u$- and $r$-band imaging surveys employed by UNIONS (gwyn_unions_2025). For the VLASS catalog, by comparing the Epoch 1 catalog with Gaia DR2, the VLASS component positional accuracy is $\sim 0.5\arcsec$ at $\mathrm{DEC.>-20\ deg^{2}}$. Therefore, the current standard deviations of the positional errors for our UNIONS-VLASS sources are reasonable compared to the entire VLASS survey, suggesting that the astrometric offsets between optical and radio pairs are attributed to the intrinsic uncertainties of VLASS QL images.
In VLASS Epoch 1, the positional error peaks at $0.0\arcsec$ in R.A. and peaks at $0.25\arcsec$ in DEC. Our UNIONS-VLASS sources based on Epoch 2 observations show a similar trend in R.A. with an improved astrometric accuracy in DEC. (${\Delta\delta}_{\mathrm{peak}}\sim 0.0\arcsec$) compared to VLASS Epoch 1. However, there is no significant improvement in the typical astrometry of $\sim 0.5\arcsec$, which is much worse than that of FIRST ($0.2\arcsec$). This is more of a problem of the rapid CLEAN procedure adopted by VLASS QL products. More accurate astrometry can be expected when Epoch 3 and epoch-concatenated data are delivered, as well as for the SE images.